whats 3(x-2)= 30.6
write this fraction as a percent 4/9
write 22.089 in fractional notation
whats the LCM 125 and 175

Answer by stano(2)

3(x-2)=30.6, simplify. 3*(x-2)=3x-6=30.6, 3x=36.6, x=36.6/3, x=12.2